Determine whether the graphs of the pair of lines are parallel.
x+9=y
y-x=-5
Answer:
The slope of line x+9=y is ? or is it undefined ?
The slope of line y-x=-5 is ? or is it undefined ?
Are the graphs of the given equations parallel?

Answer by mananth(16946)

x+9=y
write in the form of y=mx+b
y=x+9
slope = 1
y-x=-5
y=x-5
slope = 1
since the slopes are same the lines are parallel